how much pure acid should be mixed with 2 gallons of a 40% acid solution in order to get a 70% acid solution. Can you please show the formula for mixture problems. thank you.
Answer by stanbon(75887)

how much pure acid should be mixed with 2 gallons of a 40% acid solution in order to get a 70% acid solution.
-----
40% solution DATA;
amount = 2 gallons ; amount of acid = 0.40(2)=0.8 gallons
---------
100% solution DATA:
amount = x gallons ; amount of acid = 1.00x=x gallons
-------------
70% solution DATA:
amount= 2+x gallons; amount of acid = 0.7(2+x)=1.4 +0.7x gallongs
-------------
EQUATION:
acid + acid = acid
0.8 + x = 1.4+0.7x
0.3x=0.6
x=2 gallons (amount of pure acid that must be added)
Cheers,
stan H.